# Near-Optimal Very Low Rank Approximation at Sublinear Cost

### v16

@@ 1.5 Organization of our paper
We present our accelerated oversampling LRA algorithm in Sec. 2, discuss superfast computation of a crude LRA in Sec. 3, present our iterative LRA algorithm in Sec. 4, and cover our numerical experiments in Sec. 5. We devote our short Sec. 6 to conclusions. In Appendix A we describe some small families of matrices for which any superfast LRA algorithm fails. In Appendix B we compute $r$-top SVD of an LRA of a low-rank matrix superfast. In Appendix C we recall CUR decomposition and CUR LRA and transform an $r$-top SVD of a matrix into its rank-$r$ CUR LRA. In Appendix D we generate abridged SRHT matrices, whose incorporation as test matrices makes random sampling LRA superfast.
@@ 2 LRAs of $(r,\rho)$-matrices: generic algorithm and its analysis
[7, Cor. 8.6.2]. For a pair of $m\times n$ matrices $M$ and $M+E$ it holds that
@@
In the following algorithm we elaborate upon the outline of Sec. 1.2 assuming that we are given two black box subroutines that output a crude LRA of a matrix.
@@
    Compute its $r$-top SVD (cf. Appendix B), which is SVD of $X:=M(\rho)_{r}$.
@@
Upper bound (2.3) is within $2\nu$ from the lower bound $\sigma_{r+1}(M)$, but can be very pessimistic, e.g., $||E||=\sigma_{r+1}(M)$ where $M(\rho)=M_{\rho}$.
@@
[The choice of an upper rank $\rho$.] We should choose a larger value $\rho$ to decrease the norm bound $\nu$ but should keep $\rho$ small enough to perform at stage 2 fast or superfast. For a try-and-error selection among various candidate values of $\rho$, we can compare either the ratios $\frac{||M-M(\rho)||_{1}}{||M-M(r)||_{1}}$ or, at a higher computational cost, the values $\nu(M,\rho)$.
@@
[A simplified version.] In Sec. 5.3 we test a simplified version of Alg. 2.1 where we impose no upper bound $\mu$ and output the error norms $||E||=||E(\rho)||$ for $\rho=ir$ and $i=2,3,4,5$. In practice, the choice of integers $\rho$ and then $r$ should adaptive, towards minimizing the error norms of the resulting rank-$\rho$ and rank-$r$ approximations of $M$.
3 LRA by means of random sampling with sparse multipliers
The random sampling LRA algorithms of [11, 33] are close to optimal with a high probability but are not superfast under various customary choices of dense test matrices such as Gaussian, Rademacher, SRHT, and SRFT. In the case of VLRA (where $(r+p)^{2}\ll n$ in the HMT algorithm or $kl\ll n$ in the TUYC and GN algorithms) one can accelerate the computations to superfast level by using sparse test matrices, e.g., Ultra-Sparse Rademacher test matrices [33, Sec. 3.9], [3], which depend on a positive integer parameter $s$, such that a test matrices are sparse and support superfast LRA for smaller values $s$. It is a challenge to support our VLRA with these algorithms and test matrices, but in this paper we tested just TYUC algorithm with sparse abridged SRHT matrices, specified in Appendix D and also supporting superfast VLRA. For other candidate sparse test matrices see [11, Remark 4.6], [7, Sec. 5.1.8].
According to [13], such acceleration tends to make the output LRAs somewhat less reliable, and indeed in our tests LRAs computed at stage 1 of Alg. 2.1 tended to be cruder with abridged SRHT test matrices than with Gaussian random ones. Stage 2, however, tended to fix this discrepancy where the ratio $\frac{\sigma_{r+1}}{\sigma_{\rho+1}}$ was large enough.
@@ 4 Iterative refinement of LRA of an $(r,\rho)$-matrix with compression
Next we elaborate upon iterative refinement of LRA of an $(r,\rho)$-matrix, sketched in Sec. 1.3. We allow at most $k$ refinement steps for a fixed tolerance $k$. As usual for iterative refinement we perform some additions/subtractions at updating steps with a higher precision, and we increase a lower bound on the upper rank $\rho$ from $r+1$ to $2r$.
@@
as in Alg. 2.1, except that the input of Alg. 4.1 additionally includes a positive integer $k$, which bounds the number of recursive steps allowed, and an Alg. $\mathbb{S}$ (any black box subroutine) that with a higher precision adds or subtracts a pair of black box $m\times n$ matrices.
@@
    Compute SVD of $X_{i+1}:=(M_{i}(\rho))_{r}$. [This compresses $M_{i}(\rho)$.]
@@
    Otherwise output FAILURE if $i=k$.
@@
[The computational cost bound, the choice of an upper rank, and variations of Alg. 4.1.] As in Alg. 2.1 maximize the upper rank $\rho$ to decrease the output error norm, being only limited by the assumption that the overall computational cost is dominated at stage 1 (cf. (4.1)). We can allow an upper rank $\rho$ to vary as $i$ varies as long that it stays much lower than $n$ for all $i$. In Sec. 5.3 we test a variant of Alg. 4.1 where we fix $k=3$, and let $\rho:=r$ at the first iteration (for $i=0$) and $\rho:=2r$ at the next iterations.
@@
For the error matrices $E_{i}:=M-M_{i}(\rho)$ of Alg. 4.1, write $\nu_{i}:=||E_{i}||,~{}i=0,1,\dots,k-1.$ Then it holds that
@@
$\sigma_{\rho+r+1}(M)\leq\sigma_{\rho+1}(E_{i})\leq\sigma_{\rho-r+1}(M),~{}i=0,1,\dots,k-1.$ (4.3)
@@ 5.1 Input Matrices
Our third input matrix, from the discretization of a single layer potential (SLP) operator, has size $1024\times 1024$ (see [11, Sec. 7.1] for more details).
@@
We also computed rank-$r$ approximations for the three matrices called StreamVel, MinTemp, and MaxCut and used in numerical experiments in [34, Sec. 7.3.2].
@@
For both input classes we fixed target rank $r=20$ and set it equal to 10 for the following three classes of synthetic $1024\times 1024$ matrices, generated according to the recipes in [34, Sec. 7.3.1], which involved effective rank parameter $R=20$.
@@
(i) Fast (although not superfast) randomized algorithm of [33] with Gaussian random multipliers $F$ of size $(2\rho)\times m$ and $H$ of size $n\times\rho$.
(ii) Its accelerated superfast randomized variant where abridged (length 3) SRHT multipliers $F\in\mathbb{C}^{(2\rho)\times 1024}$ and $H\in\mathbb{C}^{1024\times\rho}$ replaced Gaussian ones.
@@
For both algorithms we repeated the tests 100 times per test set with independent random choices of multipliers $F$ and $H$, that is, per the triple $\{r,\rho,{\rm TYPE}\}$ for Alg. 2.1 or the pair $\{r,{\rm TYPE}\}$ for Alg. 4.1 where TYPE stands for an input matrix type. We, however, ran just 10 tests for the StreamVel and MinTemp matrices, which have significantly larger sizes.
In Tables 5.1, 5.2, 5.5, and 5.6 we display the output error norms in our tests with both Gaussian random and abridged SRHT test matrices, which were more or less in the same range for both matrix types. All our tests with the synthetic input matrices from [34, Sec. 7.3.1] wit abridged SRHT test matrices have failed, that is, have output large error norms. In Tables 5.3 and 5.4 we display the results of these tests but only with Gaussian random test matrices, where the output LRAs were quite accurate, except for LRAs of high noise and slow decay matrices.
Alg. 2.1 failed for the matrix Shaw and target rank $r=20$ with both Gaussian and abridged SRHT test matrices. This could have been predicted because the spectrum of the singular values $\sigma_{j}(M)$ of these matrices $M$ is flat for $j>20$ (decreasing roughly from $10^{-13}$ to between $10^{-15}$ and $10^{-16}$); in Table 5.1 we display the test results for the matrix Shaw and target rank $r=19$, which are about as favorable as for Gravity and Slow Decay.
@@
Input Matrix Multiplier $\rho$
@@
Input Matrix Multiplier 1st Itr. 2nd Itr. 3rd Itr.
@@
Table 5.3: Test results for Alg. 2.1 on the synthetic input matrices from [34, Sec. 7.3.1] using Gaussian test matrices.
@@
Table 5.4: Test results for Alg. 4.1 on the synthetic input matrices from [34, Sec. 7.3.1] using Gaussian multipliers.
Input Multiplier $\rho$
@@
Table 5.5: Test results for Alg. 2.1 on the real-world input matrices from [34, Sec. 7.3.2]
Input Matrix Multiplier Type 1st Itr. 2nd Itr. 3rd Itr.
@@
Table 5.6: Test results for Alg. 4.1 for $\rho=2r$ on real-world input matrices from [34, Sec. 7.3.2].
@@ 6 Conclusions
Rank-$r$ approximation of an $m\times n$ matrix $M$ is only interesting if there is a gap in the spectrum of its top (largest) $r+1$ singular values $\sigma_{j}(M)$. For the $(r,\rho)$-matrices $M$ having also a gap in the next segment of their spectra – from $(r+1)$st to $(\rho+1)$st largest singular values – randomized oversampling algorithms output quite accurate LRA.
We extended these observations towards devising superfast (sublinear cost) variation of LRA algorithms based on random sampling and oversampling with sparse test matrices and to improving accuracy of the resulting LRA in the case of $(r,\rho)$-matrices $M$ where $r<\rho\ll\sqrt{n}$. We provided some formal and experimental support for the efficiency of the resulting LRA algorithms.
There are various natural extensions of our study, which can be practically competitive. E.g., one can devise and test similar variations of the TYUC and GN algorithms with sparse test matrices such as Ultra-Sparse Rademacher test matrices [33, Sec. 3.9] or sparsified random Givens rotations by Rokhlin and Tygert (see [11, Remark 4.6]).666[26, 27, 28] generated abridged SRFT matrices similarly to abridged SRHT matrices and in extensive tests observed similar efficiency of both of these matrix classes.
Our study in Secs. 4 and 5 has showed no benefits for using iterative refinement of LRA versus oversampling so far, but one can keep comparing their efficiency.
@@
Appendix A Small families of matrices that are hard for superfast LRA
Any superfast algorithm fails to compute a close LRA of the following small families of matrices.
@@ Appendix A Small families of matrices that are hard for superfast LRA
Let $\Delta_{i,j}$ denote an $m\times n$ matrix of rank 1 filled with 0s except for its $(i,j)$th entry filled with 1. The $mn$ such matrices $\{\Delta_{i,j}\}_{i,j=1}^{m,n}$ form a family of $\delta$-matrices. We also include the $m\times n$ null matrix $O_{m,n}$ filled with 0s into this family. Now fix any superfast algorithm; it does not access the $(i,j)$th entry of its input matrices for some pair of $i$ and $j$. Therefore it outputs the same approximation of the matrices $\Delta_{i,j}$ and $O_{m,n}$, with an undetected error at least 1/2. Arrive at the same conclusion by applying the same argument to the set of $mn+1$ small-norm perturbations of the matrices of the above family and to the $mn+1$ sums of the latter matrices with any fixed $m\times n$ matrix of low rank. Likewise, we can verify that any superfast randomized LRA algorithms fail fails on those matrix families with error probability not close to 0.
@@ Appendix B Computation of $r$-top SVD of a matrix product
The algorithm involves $(m+n+k)k$ scalars and $O((m+n)k^{2})$ flops (cf. [7, Figure 8.6.1]), and so it is superfast if $k^{2}\ll n\leq m$. Its correctness follows from equations $AB=U_{A}WV^{*}_{B}$, $W=U_{W}\Sigma_{W}V^{*}_{W}$, and $\Sigma_{W}=\operatorname{diag}(\Sigma,\Sigma^{\prime})$. One can readily extend this algorithm to computing $r$-top SVD of a matrix product $ABC$ by using $O(mk^{2}+n\ell^{2})$ flops and $(m+k)k+(n+\ell)\ell$ scalars where $A\in\mathbb{R}^{m\times k}$, $B\in\mathbb{R}^{k\times\ell}$, and $C\in\mathbb{R}^{\ell\times n}$.
Appendix C CUR decomposition, CUR LRA, and its link to top SVD
C.1 CUR decomposition and CUR approximation
Next we briefly revisit CUR LRA, which is particularly memory efficient. For five integers $k$, $\ell$, $m$, $n$, and $r$ satisfying (1.2), an $m\times n$ matrix $M$, and two sets $\mathcal{I}\subseteq\{1,\dots,m\}$ and $\mathcal{J}\subseteq\{1,\dots,n\}$ of cardinalities $k$ and $\ell$, respectively, define the submatrices
$R:=M_{\mathcal{I},:}:=(m_{i,j})_{i\in\mathcal{I};j=1,\dots,n},C:=M_{:,\mathcal{J}}:=(m_{i,j})_{i=1,\dots,m;j\in\mathcal{J}},~{}{\rm and}~{}G:=M_{\mathcal{I},\mathcal{J}}:=(m_{i,j})_{i\in\mathcal{I};j\in\mathcal{J}}$
of $M$ and its CUR approximation of rank at most $r$ given by the matrix $CUR$ with its generator $G$ and nucleus $U$, set to be $(G_{r})^{+}$, the Moore–Penrose pseudo inverse of $G_{r}$, although we recall notation $G$ for nucleus and CGR for the factorization in [21], originated with the decomposition itself in the 1990s [8]. In Appendix C.2 we denote nucleus with $N$ and the decomposition with CNR for convenience.
Given a $k\times\ell$ generator submatrix $G$ of $M$, we can compute CUR approximation of $M$ by using $k\ell$ scalars and $O(k\ell\min\{k,\ell\})$ flops.
Theorem C.1.
(See [28].) For matrices $C$, $U$, and $R$ defined above, $M=CUR$ if and only if $\operatorname{rank}(G)=\operatorname{rank}(M)$.
If $\operatorname{rank}(G)<\operatorname{rank}(M)$, then the norm $||CUR-M||$ can be large.
C.2 From $r$-top SVD of a matrix to its CUR/CNR decomposition
The algorithm of Appendix B transforms a CUR decomposition into a top SVD. The next algorithm makes converse transition. It uses $(m+n+r)r$ scalars and order of $(m+n)r^{2}$ flops to compute CUR decomposition having an $r\times r$ generator for a matrix given with its $r$-top SVD, for $r\leq n\leq m$. By combining this algorithm with that of Appendix B we can transform an LRA of a matrix into its CUR LRA at a low computational cost.
Algorithm C.1.
[From $r$-top SVD to CNR decomposition.] [^7] We denote nucleus by $N$ rather than $U$ or $G$ to avoid conflicts with the factor $U$ in SVD and generator $G$ of CUR of Sec. C.1.
Input:
Three integers $m$, $n$, and $r$ satisfying $r\leq n\leq m$ and four matrices $M\in\mathbb{C}^{m\times n}$, $\Sigma\in\mathbb{C}^{r\times r}$ (nonsingular diagonal), $U\in\mathbb{C}^{m\times r}$, and $V\in\mathbb{C}^{n\times r}$ (both with orthonormal columns) such that $M:=U\Sigma V^{*}$ is SVD.
Output:
Three matrices $C\in\mathbb{C}^{m\times r}$, $N\in\mathbb{C}^{r\times r}$, and $R\in\mathbb{C}^{r\times n}$ such that $C$ and $R$ are submatrices made up of $r$ columns and $r$ rows of $M$, respectively, and
$M=CNR.$
Computations:
By applying to the matrices $U$ and $V$ the algorithms of [6] or [23] compute two nonsingular $r\times r$ submatrices $U_{\mathcal{I},:}$ and $V^{*}_{:,\mathcal{J}}$.
Compute and output the CUR factors $C=U\Sigma V^{*}_{:,\mathcal{J}}$, $R=U_{\mathcal{I},:}\Sigma V^{*}$, and $N:=(V_{:,\mathcal{J}}^{*})^{-1}\Sigma^{-1}U_{\mathcal{I},:}^{-1}$.
Correctness verification. Substitute the expressions for $C$, $N$ and $R$ and obtain $CNR=(U\Sigma V^{*}_{:,\mathcal{J}})((V_{:,\mathcal{J}}^{*})^{-1}\Sigma^{-1}U_{\mathcal{I},:}^{-1})(U_{\mathcal{I},:}\Sigma V^{*})=U\Sigma V^{*}$.
Cost bounds. The algorithm uses $O(mr^{2})$ flops and $(m+n+r)r$ scalars; these cost bounds are sublinear where $r^{2}\ll n$ and are dominated at stage 2.
Let us also estimate the norm of the nucleus $||N||$. Observe that $\operatorname{rank}(V_{:,\mathcal{J}})=\operatorname{rank}(\Sigma)=\operatorname{rank}(U_{\mathcal{I},:})=r$ and $||N||\leq||V_{:,\mathcal{J}}^{-1}||~{}||\Sigma^{-1}||~{}||U_{\mathcal{I},:}^{-1}||$ and recall that $||\Sigma^{-1}||=||M^{+}||=1/\sigma_{r}(M)$.
Write $t_{q,s,h}^{2}:=(q-s)sh^{2}+1$, allow any choice of $h>1$, say, $h=1.1$, and then recall that $||U_{\mathcal{I},:}^{+}||\leq t_{m,k,h}^{a}$, $||(V_{:,\mathcal{J}}^{*})^{+}||\leq t_{n,\ell,h}^{a}$, and consequently
$||N||\leq t_{m,r,h}^{a}t_{n,r,h}^{a}/\sigma_{r}(M)$
where $a=1$ if we apply the algorithms of [6] at stage and $a=2$ if we apply those of [23].
Appendix D Generation of abridged SRHT matrices

### v17

@@ 1.7 Organization of our paper
We present our accelerated oversampling LRA algorithm in Sec. 2, discuss fast computation of a crude LRA in Sec. 3, present our iterative LRA algorithm in Sec. 4, and cover our numerical experiments in Sec. 5. We devote our short Sec. 6 to conclusions. In Appendix A we describe some small families of matrices for which any LRA algorithm fails unless it access all entries of an input matrix. In Appendix B we fast compute $r$-top SVD of an LRA of a low-rank matrix. In Appendix C we generate abridged SRHT matrices, whose incorporation as sketch matrices makes random sampling LRA run faster.
@@ 2 LRAs of $(r,\rho)$-matrices: generic algorithm and its analysis
[6, Cor. 8.6.2]. For a pair of $m\times n$ matrices $M$ and $M+E$ it holds that
@@
In the following algorithm we elaborate upon the outline of Sec. 1.3 assuming that we are given black box subroutines that output a crude LRA and an estimate for a spectral norm of a matrix.
@@
    Compute its $r$-top SVD (cf. [19, Appendix B]), which is SVD of $X:=M(\rho)_{r}$.
@@
Upper bound (2.3) is within $2\nu$ from the lower bound $\sigma_{r+1}(M)$ but can be quite pessimistic, e.g., $||E||=\sigma_{r+1}(M)$ where $M(\rho)=M_{\rho}$.
@@
[The choice of an upper rank $\rho$.] We should choose a larger value $\rho$ to decrease the norm bound $\nu$ but should choose it smaller to perform stage 2 faster. For a try-and-error selection among various candidate values of $\rho$, we can compare either the ratios $\frac{||M-M(\rho)||_{1}}{||M-M(r)||_{1}}$ or, at a higher computational cost, the values $\nu(M,\rho)$.
@@
[A simplified version.] In Sec. 5.3 we test a simplified version of Alg. 2.1 where we impose no upper bound $\mu$ and output the error norms $||E||=||E(\rho)||$ for $\rho=ir$ and $i=2,3,4,5$. In practice, the choice of integers $\rho$ and then $r$ should be adapted towards minimizing the error norms of the computed rank-$\rho$ and rank-$r$ approximations of $M$.
3 LRA by means of random sampling with sparse sketch matrices
The random sampling LRA algorithms of [10, 27] are close to optimal with a high probability under various customary choices of dense sketch matrices such as Gaussian, Rademacher, SRHT, and SRFT. In the case of VLRA (where $(r+p)^{2}\ll n$ in the HMT algorithm or $kl\ll n$ in the TUYC and GN algorithms) one can accelerate the computations by using sparse sketch matrices, such as Ultra-Sparse Rademacher sketch matrices [27, Sec. 3.9], [2], which depend on a positive integer parameter $s$, such that sketch matrices are sparse and support faster LRA for smaller values $s$. In the present paper we tested just TYUC algorithm with sparse abridged SRHT matrices, specified in Appendix C and also supporting faster VLRA. For other candidate sparse sketch matrices see [10, Remark 4.6], [6, Sec. 5.1.8].
According to [11], such acceleration tends to make the output LRAs somewhat less reliable, and in our tests LRAs computed at stage 1 of Alg. 2.1 tended to be cruder with abridged SRHT sketch matrices than with Gaussian random ones. Stage 2, however, tended to fix this discrepancy where the ratio $\frac{\sigma_{r+1}}{\sigma_{\rho+1}}$ was large enough.
@@ 4 Iterative refinement of LRA of an $(r,\rho)$-matrix with compression
Next we elaborate upon iterative refinement of LRA of an $(r,\rho)$-matrix, sketched in Sec. 1.5. As a basis we apply Alg. 2.1, but other efficient LRA algorithms can be applied instead of it. We allow at most $h$ refinement steps for a fixed tolerance $h$. As usual for iterative refinement we perform some additions/subtractions at updating steps with a higher precision, and we increase a lower bound on the upper rank $\rho$ from $r+1$ to $2r$.
@@
as in Alg. 2.1, except that the input of Alg. 4.1 additionally includes a positive integer $h$, which bounds the number of recursive steps allowed, and an Alg. $\mathbb{S}$ (any black box subroutine) that with a higher precision adds or subtracts a pair of black box $m\times n$ matrices.
@@
    Compute SVD of $X_{i+1}:=(M_{i}(\rho))_{r}$. [This enables us to compress $M_{i}(\rho)$.]
@@
    Otherwise output FAILURE if $i=h$.
@@
[The computational cost bound, the choice of an upper rank, and variations of Alg. 4.1.] As in Alg. 2.1 maximize the upper rank $\rho$ to decrease the output error norm, being only limited by the assumption that the overall computational cost is dominated at stage 1 (cf. (4.1)). We can allow an upper rank $\rho$ to vary as $i$ varies as long that it stays much lower than $n$ for all $i$. In Sec. 5.3 we test a variant of Alg. 4.1 where we fix $h=3$, and let $\rho:=r$ at the first iteration (for $i=0$) and $\rho:=2r$ at the next iterations.
@@
For the error matrices $E_{i}:=M-M_{i}(\rho)$ of Alg. 4.1, write $\nu_{i}:=||E_{i}||,~{}i=0,1,\dots,h-1.$ Then it holds that
@@
$\sigma_{\rho+r+1}(M)\leq\sigma_{\rho+1}(E_{i})\leq\sigma_{\rho-r+1}(M),~{}i=0,1,\dots,h-1.$ (4.3)
@@ 5.1 Input Matrices
Our third input matrix, from the discretization of a single layer potential (SLP) operator, has size $1024\times 1024$ (see [10, Sec. 7.1] for more details).
@@
We also computed rank-$r$ approximations for the three matrices called StreamVel, MinTemp, and MaxCut and used in numerical experiments in [28, Sec. 7.3.2].
@@
For both input classes we fixed target rank $r=20$ and set it equal to 10 for the following three classes of synthetic $1024\times 1024$ matrices, generated according to the recipes in [28, Sec. 7.3.1], which involved effective rank parameter $R=20$.
@@
(i) Fast randomized algorithm of [27] with Gaussian random sketch matrices $F$ of size $(2\rho)\times m$ and $H$ of size $n\times\rho$.
(ii) Its accelerated randomized variant where abridged (length 3) SRHT sketch matrices $H\in\mathbb{C}^{1024\times\rho}$ replaced Gaussian ones.
@@
For both algorithms we repeated the tests 100 times per test set with independent random choices of sketch matrices $F$ and $H$, that is, per the triple $\{r,\rho,{\rm TYPE}\}$ for Alg. 2.1 or the pair $\{r,{\rm TYPE}\}$ for Alg. 4.1 where TYPE stands for an input matrix type. We, however, ran just 10 tests for the StreamVel and MinTemp matrices, which have significantly larger sizes.
In Tables 5.1, 5.2, 5.5, and 5.6 we display the output error norms in our tests with both Gaussian random and abridged SRHT sketch matrices, which were more or less in the same range for both matrix types. All our tests with the synthetic input matrices from [28, Sec. 7.3.1] wit abridged SRHT sketch matrices have failed, that is, have output large error norms. In Tables 5.3 and 5.4 we display the results of these tests but only with Gaussian random sketch matrices, where the output LRAs were quite accurate, except for LRAs of high noise and slow decay matrices.
Alg. 2.1 failed for the matrix Shaw and target rank $r=20$ with both Gaussian and abridged SRHT sketch matrices. This could have been predicted because the spectrum of the singular values $\sigma_{j}(M)$ of these matrices $M$ is flat for $j>20$ (decreasing roughly from $10^{-13}$ to between $10^{-15}$ and $10^{-16}$); in Table 5.1 we display the test results for the matrix Shaw and target rank $r=19$, which are about as favorable as for Gravity and Slow Decay.
@@
Input Matrix Sketch Matrix $\rho$
@@
Input Matrix Sketch Matrix 1st Itr. 2nd Itr. 3rd Itr.
@@
Table 5.3: Test results for Alg. 2.1 on the synthetic input matrices from [28, Sec. 7.3.1] using Gaussian sketch matrices.
@@
Table 5.4: Test results for Alg. 4.1 on the synthetic input matrices from [28, Sec. 7.3.1] using Gaussian sketch matrices.
Input Sketch Matrix $\rho$
@@
Table 5.5: Test results for Alg. 2.1 on the real-world input matrices from [28, Sec. 7.3.2]
Input Matrix Sketch Matrix Type 1st Itr. 2nd Itr. 3rd Itr.
@@
Table 5.6: Test results for Alg. 4.1 for $\rho=2r$ on real-world input matrices from [28, Sec. 7.3.2].
@@ 6 Conclusions
There are various natural extensions of our study, which can be practically competitive. E.g., one can devise and test combinations of Alg. 2.1 and of the TYUC and GN algorithms with various sparse sketch matrices such as Ultra-Sparse Rademacher sketch matrices [27, Sec. 3.9], sparsified random Givens rotations by Rokhlin and Tygert (see [10, Remark 4.6]), or even further sparsified count sketch matrices in Clarkson and Woodruff [2, 3].444[20, 21, 22] generated abridged SRFT matrices similarly to abridged SRHT matrices and in extensive tests observed similar efficiency of both of these matrix classes. Instead of random sampling LRA with sparse sketch matrices one can apply Cross-Approximation algorithm for LRA (see [17] and the references therein).
Our comparative study has showed no significant benefits of applying iterative refinement to LRA versus standalone Alg. 2.1, but one can further explore this comparison.
@@
Appendix A Small families of matrices that are hard for fast LRA
Any algorithm that does not access all entries of an input matrix fails to compute a close LRA of the following small families of matrices.
@@ Appendix A Small families of matrices that are hard for fast LRA
Let $\Delta_{i,j}$ denote an $m\times n$ matrix of rank 1 filled with 0s except for its $(i,j)$th entry filled with 1. The $mn$ such matrices $\{\Delta_{i,j}\}_{i,j=1}^{m,n}$ form a family of $\delta$-matrices. We also include the $m\times n$ null matrix $O_{m,n}$ filled with 0s into this family. Now fix any algorithm that does not access the $(i,j)$th entry of its input matrices for some pair of $i$ and $j$. Such n algorithm outputs the same approximation of the matrices $\Delta_{i,j}$ and $O_{m,n}$, with an undetected error at least 1/2. We arrive at the same conclusion by applying the same argument to the set of $mn+1$ small-norm perturbations of the matrices of the above family and to the $mn+1$ sums of the latter matrices with any fixed $m\times n$ matrix of low rank. Likewise, we can verify that any randomized LRA algorithms that does not access ail entries of an input matrix fails on those matrix families with error probability not close to 0.
@@ Appendix B Computation of $r$-top SVD of a matrix product
The algorithm involves $(m+n+k)k$ scalars and $O((m+n)k^{2})$ flops (cf. [6, Figure 8.6.1]), running particularly fast where $k\ll n\leq m$. Its correctness follows from equations $AB=U_{A}WV^{*}_{B}$, $W=U_{W}\Sigma_{W}V^{*}_{W}$, and $\Sigma_{W}=\operatorname{diag}(\Sigma,\Sigma^{\prime})$. One can readily extend this algorithm to computing $r$-top SVD of a matrix product $ABC$ by using $O(mk^{2}+n\ell^{2})$ flops and $(m+k)k+(n+\ell)\ell$ scalars where $A\in\mathbb{R}^{m\times k}$, $B\in\mathbb{R}^{k\times\ell}$, and $C\in\mathbb{R}^{\ell\times n}$.
Appendix C Generation of abridged SRHT matrices